OpenStudy (anonymous):

Use the quadratic equation y = x^2 + 3x – 10. Where does the vertex of the parabola lie? A. below the x-axis B. above the x-axis C. on the y-axis D. on the x-axis

OpenStudy (anonymous):

Familiar with completing the square?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Umm how about factoring?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[y=x^2+3x-10\] Now we need to find 2 numbers that when you add them together it equals 3 and when you multiply it equals -10 5 and -2 5-2=3 5*-2=-10 \[x^2-2x+5x-10=0\] \[(x^2-2x)+(5x-10)=0\] \[x(x-2)+5(x-2)=0\] \[(x-2)(x+5)=0\] \[ x=2 \text{ and } x=-5\] Do you follow this so far?

OpenStudy (anonymous):

Yes, I understand so far :)

OpenStudy (anonymous):

its 2 and 5

OpenStudy (anonymous):